OpenStudy (anonymous):

Find the sum of each arithmetic series. the first 100 even natural numbers.

OpenStudy (anonymous):

@radar @robtobey @amistre64 @ivettef365

OpenStudy (zzr0ck3r):

2+4+6+8+.....+100 200 +198 + 196+....+102 add top tp bottom 202+202+202...+202 there are 50 of these so 50(202) =

OpenStudy (amistre64):

guass already did this one lol

OpenStudy (zzr0ck3r):

I think

OpenStudy (amistre64):

gauss did 1 to 100 ... i forgot to actually comprehend the problem

OpenStudy (anonymous):

@zzr0ck3r how did u get 202

OpenStudy (zzr0ck3r):

im adding top to bottom

OpenStudy (amistre64):

\[\sum_{n=1}^{100}~2n\] \[2\sum_{n=1}^{100}~n\] \[2\frac{100(101)}{2}\]

OpenStudy (zzr0ck3r):

2 + 4 + 6 + ........................ + 100 200 + 198 + 196 + ..............................+ 102 200+2 = 202 198+4 = 202 196+6 = 202 ...... we will have 50 of these so 50(202) = 10100

OpenStudy (anonymous):

@amistre64 what did you plug in for the variables in the sigma notation

OpenStudy (amistre64):

the numbers 1 to 100; that would count off the first 100 even numbers

OpenStudy (anonymous):

waht did u plug in for n?

OpenStudy (amistre64):

\[\sum_{n=1}^{100}~2n\] \[2\sum_{n=1}^{100}~n\] \[2(1+2+3+4+...+100)\] \[or~\frac{100(2(1)+2(100))}{2}=100(1+100)\]

OpenStudy (amistre64):

the sum of n is a well rehearsed formula: n, times the sum of the first and last terms; divided in half

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome